# Vacuum polarisation effects in impulsive fields

### v1

@@ II.3 Evaluation of the integrals
As this corresponds to a lightlike boost of the on-shell variable $p_{\mu}$, the Jacobian for the change of variable is unity. We evaluate the $P_{\scriptscriptstyle\perp}$ integral in polar coordinates. The angular integral removes linear factors of $\varepsilon_{\scriptscriptstyle\perp}P_{\scriptscriptstyle\perp}$ from the numerator of the integrand by symmetry. One then arrives at, writing $P\equiv|P_{\scriptscriptstyle\perp}|$,
@@
$\frac{2}{|{\mathsf{C}}|^{3}}\int_{0}^{\frac{|C_{\scriptscriptstyle\perp}|}{2}}\!{\mathrm{d}}z\bigg{(}1-\frac{m^{2}}{m^{2}+\tfrac{1}{4}{\mathsf{C}}^{2}-z^{2}}\bigg{)}\;,$ (28)
@@ III.2 Small angle scattering and the infra-red
The integral term $W_{\scriptscriptstyle{\text{IR}}}({\mathsf{c}})$ is clearly finite as ${\mathsf{c}}\to 0$, hence the leading IR behaviour of $W({\mathsf{c}})$ is given by the delta function. The same will hold if $\Phi$ is not compactly supported but falls asymptotically to zero sufficiently quickly. Inserting this into (29) and (21) and using the result (easily verified) that $\widetilde{M}(0,0)=0$, we obtain
@@
We have thus found a relation between one-loop amplitudes in PP-waves and plane waves which is not dissimilar to those previously found at tree level. The first difference is that the relation holds only in a restricted kinematic regime. The second difference is the change in weight from linear to quadratric in $W({\mathsf{c}})$. We note this is consistent with the worldline calculations of Tarasov and Venugopalan (2019). This also seems consistent with the interpretation in Adamo et al. (2021a) that the PP-wave appears as a stochastic plane wave to the massive particles crossing it; there was a single particle in the tree-level calculations, which required one averaging of $W({\mathsf{c}})$, and while there are two particles here, in the loop, momentum conservation means they receive equal and opposite kicks from the PP-wave (represented by $W({\mathsf{c}})$ for the electron and $W^{\dagger}({\mathsf{c}})$ for the positron), hence there is only one average over $|W({\mathsf{c}})|^{2}$.
@@ III.3 Wide angle scattering and quantum reflection
Quantum reflection is the back-scattering of photons on spatially inhomogenous electromagnetic fields Gies et al. (2013, 2015). In our class of fields, exact back-scattering is typically forbidden, due to the conservation of the null momentum $\ell_{\scriptscriptstyle+}$. To see this consider the simplest setup of a head-on collision between the PP-wave and the photon, which means setting $\ell_{\scriptscriptstyle\perp}=\ell_{\scriptscriptstyle-}=0$. Back-scattering would then require $\ell^{\prime}_{\scriptscriptstyle\perp}=\ell^{\prime}_{\scriptscriptstyle+}=0$, but this is kinematically forbidden since $\ell^{\prime}_{\scriptscriptstyle+}=\ell_{\scriptscriptstyle+}\not=0$ is conserved. Away from exact back-scattering there is no kinematic obstruction, although we expect the scattering probability to be suppressed. To investigate this we can conveniently parameterise the emitted photon momentum as
@@
The first term in large brackets is the $\mathcal{O}(\eta^{0})$ term of the expansion; there is no ${\mathsf{c}}$-dependence in that factor, which allows us to use (19) to perform the ${\mathsf{c}}$-integrals, yielding a delta function $\delta^{2}(\mathsf{v}/\eta)$ which has no suppport in the considered limit, and so this term vanishes – this is a good consistency check, as it simply says that for exact back-scattering the amplitude vanishes, as we already knew it must for kinematic reasons.
@@
Unless $\Phi$, thus $W({\mathsf{c}})$, has a strong dependence on particular spatial directions, the vector integral can only generate results proportional to $v_{j}$, but these will vanish when contracted with the alternating tensor. In this case the amplitude is suppressed to (at least) quadratic order in our expansion; this emphasises the necessity of spatial inhomogenity for close-to-back scattering.
@@ IV Conclusions
In future work it would be interesting to see if the relation (38) between PP-waves and plane waves has some natural interpretation in terms of eikonal scattering. Further topics for investigation include other one-loop processes, e.g. electron spin flip, or photon scattering without flip, which we did not calculate here, and of course higher-loop processes. It would be intriguing to see how much progress could be made here, in particular in the plane wave case, given current interest in higher loop corrections in strong-field QED Fedotov (2017); Fedotov et al. (2023), and their contrasting behaviours in different physical regimes Podszus and Di Piazza (2019); Ilderton (2019b); Karbstein (2019); Mironov et al. (2020); Dunne and Harris (2021); Mironov and Fedotov (2022); Heinzl et al. (2021); Torgrimsson (2021); Podszus et al. (2022); Torgrimsson (2023).

### v2

@@ II.3 Evaluation of the integrals
As this corresponds to a lightlike boost of the on-shell variable $p_{\mu}$, the Jacobian for the change of variable is unity. We evaluate the $P_{\scriptscriptstyle\perp}$–integral in polar coordinates. The angular integral removes linear factors of $\varepsilon_{\scriptscriptstyle\perp}P_{\scriptscriptstyle\perp}$ from the numerator of the integrand by symmetry. One then arrives at, writing $P\equiv|P_{\scriptscriptstyle\perp}|$,
@@
$\frac{2}{|{\mathsf{C}}|^{3}}\int_{0}^{\frac{|{\mathsf{C}}|}{2}}\!{\mathrm{d}}z\bigg{(}1-\frac{m^{2}}{m^{2}+\tfrac{1}{4}{\mathsf{C}}^{2}-z^{2}}\bigg{)}\;,$ (28)
@@
We remark that the weak-field limit of our results is obtained by expanding (9) in powers of $e$. Keeping only terms up to linear order gives
$W({\mathsf{c}})\to{\hat{\delta}}^{2}({\mathsf{c}})+ie\widetilde{\Phi}({\mathsf{c}})\;,$ (30)
in which $\widetilde{\Phi}$ is the Fourier transform of $\Phi$. Inserting this into (21), the $e^{0}$ term in the product of $W$ factors vanishes, using the result (easily verified) that ~(0,0)=$\widetilde{\mathcal{M}}(0,0)=0$, while the linear-in-$e$ terms cancel by symmetry. This leaves
$\begin{split}i\mathcal{M}\to e^{2}\hat{\delta}(\Delta\ell_{\scriptscriptstyle+})\int\!\hat{{\mathrm{d}}}^{2}{\mathsf{c}}\,{\widetilde{\Phi}}({\mathsf{c}}){\widetilde{\Phi}}^{\dagger}({\mathsf{c}}+\Delta{\mathsf{l}})\widetilde{\mathcal{M}}(c,\Delta\ell)\;.\end{split}$ (31)
Since $\widetilde{\mathcal{M}}$ is itself $\mathcal{O}(e^{2})$, the whole result is $\mathcal{O}(e^{4})$, i.e. the weak-field limit comes from the four-point one-loop diagram in which two legs are sourced by $\Phi$, as expected in perturbation theory.
@@ III.2 Small angle scattering and the infra-red
The integral term $W_{\scriptscriptstyle{\text{IR}}}({\mathsf{c}})$ is clearly finite as ${\mathsf{c}}\to 0$, hence the leading IR behaviour of $W({\mathsf{c}})$ is given by the delta function. The same will hold if $\Phi$ is not compactly supported but falls asymptotically to zero sufficiently quickly. Inserting this into (29) and (21) and again using the result that ~(0,0)=$\widetilde{\mathcal{M}}(0,0)=0$, we obtain
@@
We have thus found a relation between one-loop amplitudes in PP-waves and plane waves which is not dissimilar to those previously found at tree level. The first difference is that the relation holds only in a restricted kinematic regime. The second difference is the change in weight from linear to quadratic in $W({\mathsf{c}})$. We note this is consistent with the worldline calculations of Tarasov and Venugopalan (2019). This also seems consistent with the interpretation in Adamo et al. (2021a) that the PP-wave appears as a stochastic plane wave to the massive particles crossing it; there was a single particle in the tree-level calculations, which required one averaging of $W({\mathsf{c}})$, and while there are two particles here, in the loop, momentum conservation means they receive equal and opposite kicks from the PP-wave (represented by $W({\mathsf{c}})$ for the electron and $W^{\dagger}({\mathsf{c}})$ for the positron), hence there is only one average over $|W({\mathsf{c}})|^{2}$.
@@ III.3 Wide angle scattering and quantum reflection
Quantum reflection is the back-scattering of photons on spatially inhomogeneous electromagnetic fields Gies et al. (2013, 2015). In our class of fields, exact back-scattering is typically forbidden, due to the conservation of the null momentum $\ell_{\scriptscriptstyle+}$. To see this consider the simplest setup of a head-on collision between the PP-wave and the photon, which means setting $\ell_{\scriptscriptstyle\perp}=\ell_{\scriptscriptstyle-}=0$. Back-scattering would then require $\ell^{\prime}_{\scriptscriptstyle\perp}=\ell^{\prime}_{\scriptscriptstyle+}=0$, but this is kinematically forbidden since $\ell^{\prime}_{\scriptscriptstyle+}=\ell_{\scriptscriptstyle+}\not=0$ is conserved. Away from exact back-scattering there is no kinematic obstruction, although we expect the scattering probability to be suppressed. To investigate this we can conveniently parameterise the emitted photon momentum as
@@
The first term in large brackets is the $\mathcal{O}(\eta^{0})$ term of the expansion; there is no ${\mathsf{c}}$-dependence in that factor, which allows us to use (19) to perform the ${\mathsf{c}}$-integrals, yielding a delta function $\delta^{2}(\mathsf{v}/\eta)$ which has no support in the considered limit, and so this term vanishes – this is a good consistency check, as it simply says that for exact back-scattering the amplitude vanishes, as we already knew it must for kinematic reasons.
@@
Unless $\Phi$, thus $W({\mathsf{c}})$, has a strong dependence on particular spatial directions, the vector integral can only generate results proportional to $v_{j}$, but these will vanish when contracted with the alternating tensor. In this case the amplitude is suppressed to (at least) quadratic order in our expansion; this emphasises the necessity of spatial inhomogeneity for close-to-back scattering.
@@ IV Conclusions
PP-waves approximate the fields of ultra-boosted charge distributions, suggesting themselves as natural field models in high-energy accelerator scenarios. Indeed, this is the context in which they are commonly studied, at least in QCD Balitsky (2001); Gelis et al. (2010); Caron-Huot (2015). An interesting question to explore in future work is when this approximation becomes appropriate in the context of future colliders searching for strong field QED effects Del Gaudio et al. (2019); Yakimenko et al. (2019); in current analyses the dominant theoretical model is based on the ‘locally constant field approximation’ Ritus (1985) which takes the high field strength – of e.g. a dense particle bunch – to be the dominant scale, rather than the ultra-relativistic velocity of that bunch. Understanding if or when the PP-wave model is more appropriate could potentially simplify predictions and analysis for future QED experiments involving ultra-boosted particles and strong fields. Indeed, understanding this point is one of the longer-term motivations behind Adamo et al. (2021b) and this paper.
It would also be interesting to see if the relation (40) between PP-waves and plane waves has some natural interpretation in terms of eikonal scattering. Further topics for investigation include other one-loop processes, e.g. electron spin flip, or photon scattering without flip, which we did not calculate here, and of course higher-loop processes. It would be intriguing to see how much progress could be made here, in particular in the plane wave case, given current interest in higher loop corrections in strong-field QED Fedotov (2017); Fedotov et al. (2023), and their contrasting behaviours in different physical regimes Podszus and Di Piazza (2019); Ilderton (2019b); Karbstein (2019); Mironov et al. (2020); Dunne and Harris (2021); Mironov and Fedotov (2022); Heinzl et al. (2021); Torgrimsson (2021); Podszus et al. (2022); Torgrimsson (2023).